Multiple Production Department Factory Overhead Rate Method Performance Gloves, Inc. produces three sizes of sports gloves:...

  1. Multiple Production Department Factory Overhead Rate Method

    Performance Gloves, Inc. produces three sizes of sports gloves: small, medium, and large. A glove pattern is first stenciled onto leather in the Pattern Department. The stenciled patterns are then sent to the Cut and Sew Department, where the glove is cut and sewed together. Performance Gloves uses the multiple production department factory overhead rate method of allocating factory overhead costs. Its factory overhead costs were budgeted as follows:

    Pattern Department overhead $100,800
    Cut and Sew Department overhead 170,200
    Total $271,000

    The direct labor estimated for each production department was as follows:

    Pattern Department 1,800 direct labor hours
    Cut and Sew Department 2,300
    Total 4,100 direct labor hours

    Direct labor hours are used to allocate the production department overhead to the products. The direct labor hours per unit for each product for each production department were obtained from the engineering records as follows:

    Production Departments Small Glove Medium Glove Large Glove
    Pattern Department 0.05 0.06 0.07
    Cut and Sew Department 0.07 0.09 0.11
    Direct labor hours per unit 0.12 0.15 0.18

    If required, round all per unit answers to the nearest cent.

    a. Determine the two production department factory overhead rates.

    Pattern Department $per dlh
    Cut and Sew Department $per dlh

    b. Use the two production department factory overhead rates to determine the factory overhead per unit for each product.

    Small glove $ per unit
    Medium glove $ per unit
    Large glove $ per unit

Check My Work
Answer #1
a
Pattern Department 56 per dlh =100800/1800
Cut and Sew Department 74 per dlh =170200/2300
b
Small glove 7.98 per unit =(56*0.05)+(74*0.07)
Medium glove 10.02 per unit =(56*0.06)+(74*0.09)
Large glove 12.06 per unit =(56*0.07)+(74*0.11)